Grade 12th passPhysical Chemistry

33g nh3=_______mole=_______molecules...........??????????????

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Pawan Prajapati
5 Years ago
Thanks

Let the no. of moles be X. Molecular weight of ammonia =17g 17 g - 1 mole 33 g - X X = 34/17=2 moles Thus, the no. of moles on 33 g of ammonia is 2 moles.


Thanks